OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x + 3y = 2 2x - 3y = 22

OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x + 3y = 2 2x - 3y = 22

OpenStudy (anonymous):

x+3y=2 2x-3y=22 x+3y=2,x=2-3y substitute in 2x-3y=22,we get 2(2-3y)-3y=22 or 4-6y-3y=22 -9y=22-4=18 y=18/-9=-2 hence x=2-3(-2)=2+6=8 solution is (8,-2)